Below are the jersey numbers of 11 players randomly selected from a football team. Find the range, variance, and standard deviation for theg ven sample data. What do the results tell us? 83 60 6 62 82 63 47 94 89 27 24 Range (Round to one decimal place as needed.)
Answer #1

Range = Maximum observation - Minimum observation = 94 - 6 = 88

Now, T2 = 45333,-= 57.9091

Sample variance =

(45333-11 × 57.90912) 844.4898

So, Sample Standard deviation = (zi-T)2-V844.4898= 29.0601 rl